10 +vely charged particals r kept fixed on xaxis at pts. x=10cm ,20cm,30cm ,.......100cm . the first partical has charge 1x10^ -8 c,2nd 8x10^ -8 c ,3rd 27x10^ -8c & so on .the 10th partical has 1000x10^ -8. find the mag. of ele. force acting on 1c charge placed at origine ? Ans -4.95x10^5 N. plz give details ,i'm in herry.

Grade:12

1 Answers

Chetan Mandayam Nayakar
13 years ago

Dear Priya,

 

Force=∑(1/4πε0)(q1q2/r2)=(1/4πε0)∑(from n=1 to n=10)((n3*10-8)/(n/10)2)

 

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

Askiitians Expert

Chetan Mandayam Nayakar  IIT Madras